OpenStudy (anonymous):

(-12m^8+6m^7-20m^6)/2m^3 please tell me and help me work out

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{ x^a }{ x^b } = x^{a-b}\] \[-\frac{ 12m^8 }{ 2m^3 } + \frac{ 6m^7 }{ 2m^3 } - \frac{20 m^6}{ 2m^3 } = -\frac{ 12 }{ 2 }m^{8-3} +\frac{ 6 }{ 2 }m^{7-3} - \frac{ 20 }{ 2 }m^{6-3}\]\[= -6m^5 + 3m^4 - 10m^3\]

OpenStudy (anonymous):

i dont understant how to work it out or how you set it up like that

OpenStudy (anonymous):

i'm not sure how to be more explicit. try to see how i got the first part from your equation. \[\frac{ a+b}{ c } = \frac{ a }{ c } + \frac{ b }{ c }\]

OpenStudy (anonymous):

oh wow im an idiot. i get it now